OpenStudy (anonymous):

Solve for x: 3/4 x + 5/8 = 4x this question isn't multiple choice, its an example question from my text book. so if some one could show me all the steps to get the answer that would be really helpful.

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

I would just start by moving x to one sides of the equation. So I would take subtract -3/4x on both sides, leaving 5/8=4x - 3/4x What would you do next?

OpenStudy (anonymous):

subtract 3/4x from 4x and get 5/8 = 3 1/4x??

OpenStudy (anonymous):

So you can leave it as 5/8=13x/4 Now multiply both sides by 4 which leave you with..... 5/2 = 13 x Now to get x on it's own you divide both sides by 13. x = _____?

OpenStudy (anonymous):

where did you get 13 from?

OpenStudy (anonymous):

13/4 is the same as 3 1/4

OpenStudy (anonymous):

ok, idk how to divide 5/2 by 13

OpenStudy (anonymous):

 you can multiply by the reciprocal